A sample solution containing sulfate ions (SO42−) was analyzed and was found to contain 3 mol of sulfate ions (SO42−).
Calculate the equivalent of sulfate ions (SO42−) present in the solution.
1
Expert's answer
2019-10-29T06:44:39-0400

Е=n·М/z=3·96/2=144g